### v3

# VGLD: Visually-Guided Linguistic Disambiguation for Monocular Depth Scale Recovery

Abstract
Monocular depth estimation can be broadly categorized into two directions: relative depth estimation, which predicts normalized or inverse depth without absolute scale, and metric depth estimation, which aims to recover depth with real-world scale. While relative methods are flexible and data-efficient, their lack of metric scale limits their utility in downstream tasks. A promising solution is to infer absolute scale from textual descriptions. However, such language-based recovery is highly sensitive to natural language ambiguity, as the same image may be described differently across perspectives and styles. To address this, we introduce VGLD (Visually-Guided Linguistic Disambiguation), a framework that incorporates high-level visual semantics to resolve ambiguity in textual inputs. By jointly encoding both image and text, VGLD predicts a set of global linear transformation parameters that align relative depth maps with metric scale. This visually grounded disambiguation improves the stability and accuracy of scale estimation. We evaluate VGLD on representative models, including MiDaS and DepthAnything, using standard indoor (NYUv2) and outdoor (KITTI) benchmarks. Results show that VGLD significantly mitigates scale estimation bias caused by inconsistent or ambiguous language, achieving robust and accurate metric predictions. Moreover, when trained on multiple datasets, VGLD functions as a universal and lightweight alignment module, maintaining strong performance even in zero-shot settings. Code will be released upon acceptance.
Figure 1: As observed in the figure above, a single image can have multiple different descriptions, and these varying descriptions can significantly affect depth estimation. In particular, the orange bounding boxes in the depth estimation maps highlight this issue, especially for the RSA method, where two semantically similar text descriptions result in substantial differences in depth estimation. In contrast, VGLD(ours) demonstrates relatively stable performance across different descriptions.
Introduction
Monocular depth estimation is a fundamental and long-standing task in computer vision, with applications ranging from autonomous driving(Schön et al. 2021), augmented reality(Ganj et al. 2023) to 3D reconstruction(Mescheder et al. 2019). The goal is to predict dense depth maps from single RGB images. However, reconstructing 3D geometry from a single image is an ill-posed problem because perspective projection causes a loss of depth dimension: any point along a projection ray corresponds to the same image coordinate. Consequently, the absolute distance from the camera to the scene cannot be directly recovered from a single view. Without camera calibration, additional sensors (e.g., IMU(Wofk et al. 2023), LiDAR(Lin et al. 2024)), or strong priors such as pre-trained depth models, scale ambiguity arises. While stereo or multi-view images can resolve scale by localizing points in 3D space, modern monocular depth estimation models are often trained on diverse datasets with varying data types and distributions—including single RGB images, video streams, and images with or without calibration parameters. These differences exacerbate the challenge of scale ambiguity, especially when deploying models across domains such as indoor and outdoor scenes.
To address scale ambiguity in monocular depth estimation, one line of work trains on multi-domain datasets (e.g., indoor and outdoor) to learn depth from domain-specific distributions(Ranftl et al. 2020; Reiner et al. 2023; Yang et al. 2024a, b). However, dataset biases limit generalization(Piccinelli et al. 2024). An alternative strategy is to leverage complementary cues shared across domains. Recent approaches explore language as a modality to resolve scale ambiguity without requiring expensive sensors (e.g., LiDAR). RSA(Zeng et al. 2024b) pioneers this direction by hypothesizing that textual descriptions can guide scale estimation and demonstrates that scale-less relative depth can be mapped to metric predictions via a language-guided global transformation.
Nevertheless, linguistic inputs are inherently ambiguous—semantically similar captions may produce inconsistent scales (see Figure 1), affecting stability. Still, language is robust to visual challenges like lighting or occlusion.
To reduce linguistic ambiguity, we propose a Visually-Guided Linguistic Disambiguation (VGLD) framework, which enriches textual inputs with semantic features extracted from the corresponding image using a CLIP Image Encoder(Radford et al. 2021). Additionally, to handle cross-domain depth variation, we introduce a Domain Router Mechanism (DRM) inspired by ZoeDepth(Bhat et al. 2023), which routes inputs to domain-specific heads for consistent metric predictions. To further stabilize training, we formulate depth scale recovery as a scalar regression task and supervise it using pseudo-labels $(k_{lm},b_{lm})$ obtained via the Levenberg-Marquardt algorithm. This nonlinear optimization technique helps guide the model toward an accurate training trajectory, enhancing robust scale recovery.
Our contributions are as follows:
  • We integrate high-level semantic information from the corresponding image alongside the textual description, thereby stabilizing the output of the scalars parameters;
  • We introduce the Domain Router Mechanism, which aids in solving the cross-domain estimation problem;
  • We leverage the Levenberg-Marquardt algorithm to optimize the training trajectory and guide the model’s training process;
  • Extensive experiments demonstrate the effectiveness of our method in both indoor and outdoor scenarios, highlighting its robustness to textual variations and strong zero-shot generalization.
Figure 2: Overview. We infer the scale $\hat{k}$ and shift $\hat{b}$ from the linguistic description and the corresponding image to transform the relative depth from the depth model into a metric depth (absolute depth in meters) prediction.
Related Work
Monocular Depth Estimation
Monocular Depth Estimation (MDE) is a fundamental task in computer vision, with its development generally following two main directions: relative depth estimation and metric depth estimation. The goal of metric depth estimation is to predict pixel-wise depth values in metric units (e.g., meters), and models are typically trained by minimizing the discrepancy between predicted and ground-truth depth maps. In contrast, relative depth estimation focuses on inferring the ordinal relationships between pixel pairs, without providing any information about scale or units. A notable early milestone in this field was Eigen et al.(Eigen, Puhrsch, and Fergus 2014), the first to apply Convolutional Neural Networks (CNNs) to MDE. More recent methods such as AdaBins (Bhat, Alhashim, and Wonka 2021), LocalBins(Bhat, Alhashim, and Wonka 2022) and Binsformer(Li et al. 2024) reformulate the depth regression problem as a classification task through depth discretization. Multi-task learning strategies have also been explored: GeoNet(Qi et al. 2018) integrates surface normal estimation, while AiT(Ning et al. 2023) incorporates instance segmentation, both to enhance depth prediction through joint training. MiDaS(Ranftl et al. 2020; Reiner et al. 2023) and Diversedepth(Yin et al. 2020) advances relative depth estimation by pretraining on a diverse mixture of datasets, achieving strong generalization across domains. In addition, diffusion-based(Viola et al. 2024; Zhang et al. 2024; Song et al. 2025) methods, such as DDP (Ji et al. 2023), Marigold (Ke et al. 2024), and GeoWizard (Fu et al. 2024), adapt powerful diffusion priors to the depth estimation task via fine-tuning, enabling significant performance gains.
Metric Depth Scale Recovery
Relative depth estimation models have emerged as strong backbones for many metric depth Scale Recovery tasks, owing to their impressive cross-domain generalization and robustness. Building on MiDaS(Ranftl et al. 2020), DPT(Ranftl, Bochkovskiy, and Koltun 2021) replaces the convolutional backbone with a Vision Transformer and adapts it to metric depth via fine-tuning on scale-annotated datasets. ZoeDepth(Bhat et al. 2023) further enhances this pipeline by introducing a powerful decoder with a metric bins module, enabling effective scale recovery through supervised fine-tuning. Depth Anything extends ZoeDepth(Bhat et al. 2023) by replacing the MiDaS(Ranftl et al. 2020) encoder with its own architecture, achieving implicit conversion from relative to metric depth.
Other methods like Metric3D(Hu et al. 2024a; Yin et al. 2023), zeroDepth(Guizilini et al. 2023) and UniDepth(Piccinelli et al. 2024) recover scale by leveraging or predicting camera intrinsics, while PromptDA(Lin et al. 2024) introduces a lightweight LiDAR prompt to guide metric estimation. RSA(Zeng et al. 2024b) proposes an alternative paradigm by aligning relative depth with metric scale using textual descriptions, enabling generalization without requiring ground-truth depth at inference. However, RSA(Zeng et al. 2024b) is sensitive to linguistic variations, where semantically similar but differently worded inputs may cause inconsistent predictions. In contrast, VGLD leverages visual semantics to guide linguistic disambiguation, enabling more robust and reliable scale recovery. By grounding ambiguous textual inputs in high-level visual context, it mitigates sensitivity to language variation and achieves consistent metric depth estimation across domains.
Language Modality for Metric Depth Estimation
Recent advances in vision-language models(Li et al. 2022; Radford et al. 2021; Jia et al. 2022), driven by large-scale pretraining, have enabled strong cross-modal representations and inspired new approaches in monocular depth estimation. DepthCLIP(Zhang et al. 2022) first applied CLIP(Radford et al. 2021) to this task by reformulating depth regression as distance classification using natural language descriptions such as ”This object is giant, close…far…”, enabling zero-shot depth prediction via CLIP’s semantic priors. Subsequent works improved adaptability in various ways: Auty et al.(Auty and Mikolajczyk 2023) introduced learnable prompts to replace fixed text tokens; Hu et al.(Hu et al. 2024b) employed codebooks to address domain shifts; and CLIP2Depth(Kim and Lee 2024) proposed mirror embeddings to eliminate reliance on explicit textual input. Other approaches such as VPD(Zhao et al. 2023) , TADP(Kondapaneni et al. 2024) , EVP(Lavreniuk et al. 2023) and GeoWizard(Fu et al. 2024) extract semantic priors from pretrained text-to-image diffusion models to support depth prediction.
Recently, Wordepth(Zeng et al. 2024a) modeled language as a variational prior by explicitly encoding object attributes (e.g., size, position) to align relative predictions with metric depth. RSA(Zeng et al. 2024b) introduced a direct constraint to recover metric scale from text, but suffers from sensitivity to linguistic variation. In contrast, VGLD combines CLIP-based visual semantics with textual input, offering more stable and robust scale predictions compared to purely language-based methods.
Figure 3: Sensitivity to variations in linguistic descriptions on the NYUv2 dataset. We focus on the estimation results under three different textual inputs (text1-3). As shown in the depth maps, the RSA method exhibits noticeable sensitivity to textual variations, leading to inconsistent predictions—particularly in the regions highlighted by orange boxes. In contrast, our proposed VGLD produces more stable and consistent depth estimates across different descriptions. Warmer colors (red) indicate closer distances, while cooler colors (blue) indicate farther distances.
[…]
Abs Rel $\downarrow$ RMSE $\downarrow$ D1 $\uparrow$ Abs Rel $\downarrow$ RMSE $\downarrow$ D1 $\uparrow$
ZoeDepth(Bhat et al. 2023) robust depth estimation $\ddagger$ 0.077 0.277 0.953 0.054 2.281 0.971
ZeroDepth(Guizilini et al. 2023) 0.074 0.269 0.954 0.053 2.087 0.968
Metric3Dv2(Hu et al. 2024a) 0.047 0.183 0.989 0.044 1.985 0.985
MiDas-1(Reiner et al. 2023) Least Squares 0.121 0.388 0.866 0.333 6.901 0.408
Levenberg Marquardt 0.056 0.218 0.969 0.091 3.373 0.925
RSA-N/K(Zeng et al. 2024b) 0.171 0.569 0.731 0.163 4.082 0.798
RSA-NK(Zeng et al. 2024b) 0.168 0.561 0.737 0.160 4.232 0.782
VGLD-N/K-T (Ours) 0.158 0.529 0.758 0.133 3.755 0.854
VGLD-N/K-I (Ours) 0.121 0.423 0.860 0.120 3.668 0.868
VGLD-N/K-TCI (Ours) 0.119 0.414 0.867 0.120 3.598 0.871
VGLD-NK-T (Ours) 0.159 0.526 0.751 0.130 3.744 0.844
VGLD-NK-I (Ours) 0.123 0.426 0.855 0.122 3.574 0.868
VGLD-NK-TCI (Ours) 0.120 0.414 0.863 0.120 3.559 0.874
MiDas-2(Ranftl et al. 2020) Least Squares 0.130 0.421 0.845 0.336 6.925 0.421
Levenberg Marquardt 0.094 0.330 0.916 0.155 4.190 0.809
VGLD-N/K-T (Ours) 0.180 0.596 0.688 0.194 5.030 0.709
VGLD-N/K-I (Ours) 0.154 0.524 0.775 0.183 4.842 0.942
VGLD-N/K-TCI (Ours) 0.151 0.507 0.789 0.178 4.806 0.748
VGLD-NK-T (Ours) 0.182 0.615 0.682 0.191 4.994 0.723
VGLD-NK-I (Ours) 0.155 0.520 0.776 0.184 4.808 0.740
VGLD-NK-TCI (Ours) 0.151 0.513 0.780 0.180 4.804 0.737
DAV2-vits(Yang et al. 2024b) Least Squares 0.122 0.392 0.866 0.330 6.737 0.423
Levenberg Marquardt 0.052 0.209 0.969 0.103 3.277 0.919
VGLD-N/K-T (Ours) 0.163 0.546 0.713 0.166 4.189 0.756
VGLD-N/K-I (Ours) 0.128 0.433 0.830 0.154 4.219 0.756
VGLD-N/K-TCI (Ours) 0.125 0.423 0.842 0.152 3.872 0.779
VGLD-NK-T (Ours) 0.161 0.539 0.714 0.164 4.287 0.752
VGLD-NK-I (Ours) 0.127 0.436 0.835 0.160 4.031 0.761
VGLD-NK-TCI (Ours) 0.127 0.434 0.835 0.153 3.980 0.772
DAV1-vits(Yang et al. 2024a) Least Squares 0.121 0.397 0.863 0.331 6.772 0.423
Levenberg Marquardt 0.057 0.230 0.967 0.112 3.375 0.897
RSA-N/K(Zeng et al. 2024b) 0.147 0.484 0.775 0.160 4.437 0.780
RSA-NK(Zeng et al. 2024b) 0.148 0.498 0.776 0.158 4.457 0.786
VGLD-N/K-T (Ours) 0.145 0.496 0.792 0.151 4.354 0.773
VGLD-N/K-I (Ours) 0.115 0.405 0.872 0.144 4.074 0.790
VGLD-N/K-TCI (Ours) 0.112 0.390 0.887 0.140 4.081 0.807
VGLD-NK-T (Ours) 0.142 0.483 0.787 0.148 4.293 0.781
VGLD-NK-I (Ours) 0.114 0.404 0.880 0.142 4.151 0.814
VGLD-NK-TCI (Ours) 0.112 0.392 0.883 0.136 4.008 0.816
Table 1: Quantitative Depth Comparison on the NYUV2 and KITTI Dataset. $\dagger$ In the Model column, MiDas-1 denotes Midas-V3.1-dpt_swin2_large_384, MiDas-2 denotes Midas-V3.0-dpt_large_384, DAV2-vits denotes Depth-Anything-V2-Small, and DAV1-vits denotes Depth-Anything-V1-Small. $\ddagger$ denotes the results of certain state-of-the-art (SOTA) absolute scale estimation models. $\ast$ In the Method column, “N” and “K” indicate models trained on the NYUv2 and KITTI datasets, respectively. For example, VGLD-N/K-TCI refers to VGLD-N-TCI when evaluated on NYUv2, and VGLD-K-TCI when evaluated on KITTI. Best results are in bold, second best are underlined.
Method
Preliminaries
The objective of monocular depth estimation is to predict continuous per-pixel depth values from a single RGB image(Eigen, Puhrsch, and Fergus 2014). We consider a dataset $\mathcal{D}=\{(I^{(n)},t^{(n)},d^{(n)}_{gt},dm^{(n)}_{gt})\}_{n=1}^{N}$ consisting of $N$ samples, where each sample includes an RGB image $I\in\mathbb{R}^{3\times H\times W}$, a corresponding linguistic description $t$, a ground-truth metric depth map $d_{gt}\in\mathbb{R}^{H\times W}$ and a ground-truth domain labels $dm_{gt}\in\{0,1\}$ which represent indoor or outdoor scene. We build upon a pretrained monocular relative depth estimation model $h_{\theta}$, which serves as the foundation for our metric depth scale recovery framework. Given an RGB image, the model predicts an inverse relative depth map $x\in\mathbb{R}^{H\times W}$, which lacks absolute scale information. To recover metric-scale depth from this scaleless prediction, we apply a global linear transformation informed by both the linguistic description and high-level visual semantics of the image. Specifically, similar to RSA(Zeng et al. 2024b), we predict a pair of scalars $(\hat{k},\hat{b})\in\mathbb{R}^{2}$ that represent the scale and shift parameters of the transformation. The final metric depth prediction is then computed as:
$\hat{d}_{pred}=\frac{1}{\hat{k}\cdot x+\hat{b}}\text{ ,where }\hat{d}_{pred}\in\mathbb{R}^{H\times W}$ (1)
VGLD
To model the relationship between the linear transformation parameters and the semantic content of both the image and its linguistic description, we leverage the CLIP model as a feature extractor. Benefiting from large-scale contrastive pretraining(Radford et al. 2021), CLIP provides a shared latent space that is well-suited for aligning object-centric visual and linguistic representations. Given an input sample $\{I,t\}$, we first extract visual and text embeddings using the CLIP image encoder and CLIP text encoder, respectively. The resulting embeddings are concatenated to form a fused representation, which is subsequently passed through a lightweight encoder network, GlobalNet—a three-layer MLP—to produce a compact 256-dimensional latent embedding used for downstream scale parameter regression.
Following ZoeDepth(Bhat et al. 2023), we employ a lightweight MLP-based classifier, referred to as the Domain Routing Mechanism (DRM), to predict the domain of the input image based on its latent embedding. We consider two domains: indoor and outdoor. The predicted domain is then used to route the latent embedding to the corresponding domain-specific scalars prediction head.
Loss Function
As illustrated in Figure 2, the VGLD model freezes the weights of both the CLIP backbone and the relative depth estimator during training, and updates only the parameters of the GlobalNet and DRM modules. These modules are jointly optimized under a unified loss function. Since VGLD focuses on predicting a pair of global scalars rather than pixel-wise metric depth values, we do not adopt the Scale-Invariant Logarithmic Loss, which is more suitable for dense depth estimation tasks. Instead, following RSA(Zeng et al. 2024b), we adopt the L1 loss, which provides a more direct and interpretable supervision signal for scalars regression. The $\mathcal{L}_{metric}$ is formulated as:
$\mathcal{L}_{metric}=\frac{1}{M}\sum_{(i,j)\in\Omega}m(i,j)\times|\hat{d}_{pred}(i,j)-d_{gt}(i,j)|,$ (2)
where $\hat{d}_{pred}$ denotes the predicted metric depth, $(i,j)\in\Omega$ represents the image coordinates, $m(\cdot)\in\{0,1\}$ denotes the binary mask map and $M$ represents the number of pixels with valid ground truth values.
To ensure correct routing to the domain-specific scalars prediction head, We introduce a domain classification loss, denoted as $\mathcal{L}_{\text{dm}}$, implemented using the cross-entropy loss:
$\mathcal{L}_{domain}=CrossEntropy(\hat{dm}_{pred},dm_{gt})$ (3)
where $\hat{dm}_{pred}\in\{0,1\}$ is the predicted domain label, and $dm_{gt}\in\{0,1\}$ is the corresponding ground-truth domain.
To guide the model towards the optimal solution, we employ an MSE loss to provide LM loss (scalars supervision) for the modules:
$\mathcal{L}_{lm}=10\times(\hat{k}-k_{\text{lm}})^{2}+(\hat{b}-b_{\text{lm}})^{2}$ (4)
where $(\hat{k},\hat{b})$ are the predicted LM scalars from VGLD, and $(k_{\text{lm}},b_{\text{lm}})$ are the corresponding pseudo-labels provided by the Levenberg-Marquardt algorithm. We assign a higher weight (10x) to the scale term $k_{\text{lm}}$ because empirical observations show that the model is more sensitive to errors in scale prediction than in shift. This design choice helps stabilize training and ensures more accurate depth scaling.
The total loss is defined as follows:
$\mathcal{L}_{total}=\mathcal{L}_{metric}+\alpha\times\mathcal{L}_{domain}+\beta\times\mathcal{L}_{lm}$ (5)
In our experiments, we set $\alpha$ and $\beta$ to 0.1, as is customary.
Figure 4: Sensitivity to variations in linguistic descriptions on the KITTI dataset. Similar to Figure 3, we focus on the differences within the orange boxes across the three textual inputs. Note that we use LM fitting results instead of the ground-truth depth map for visualization, as the KITTI ground-truth data is too sparse to yield meaningful visual comparisons. Warmer colors (red) indicate closer distances, while cooler colors (blue) indicate farther distances.
Experiments
Experimental Settings
Dataset. We primarily train on two datasets: NYUv2(Silberman et al. 2012) and KITTI(Geiger, Lenz, and Urtasun 2012), representing indoor and outdoor scenes, respectively. NYUv2 contains images with a resolution of 480x640, with depth values ranging from 0 to 10 meters. In accordance with the official dataset split(Lee et al. 2019), we use 24,231 image-depth pairs for training and 654 image-depth pairs for testing. KITTI is an outdoor dataset collected from equipment mounted on a moving vehicle, with depth values ranging from 0 to 80 meters. Following KBCrop(Uhrig et al. 2017), all RGB images and depth maps are cropped to a resolution of 1216x352. We adopt the Eigen split(Eigen, Puhrsch, and Fergus 2014), which includes 23,158 training images and 652 test images, to train and evaluate our method. Additionally, we report zero-shot generalization results on SUNRGBD(Song et al. 2015), which includes 5,050 test images, DIML Indoor(Cho et al. 2021), which contains 503 validation images and DDAD(Guizilini et al. 2020), which contains 3950 validation images.
Relative Depth Models. We use MiDaS 3.1(Reiner et al. 2023) with the dpt_swin2_large_384 model (213M parameters), MiDaS 3.0(Ranftl et al. 2020) with the dpt_large_384 model (123M parameters), DepthAnything(Yang et al. 2024a) with DepthAnything-Small model (24.8M parameters), and DepthAnything v2(Yang et al. 2024b) with DepthAnything-V2-Small model (24.8M parameters).
The Proposed Models. For clarity, we denote the proposed models as VGLD-{dataset}-{method}. The {dataset} refers to the training datasets, which include ”N” for NYUv2, ”K” for KITTI, and ”NK” for both NYUv2 and KITTI. The {method} refers to the type of embeddings used: ”T” for text embeddings only, ”I” for visual embeddings only, and ”TCI” for both text and visual embeddings (i.e., Fusion Embeddings, as shown in Figure 2).
Evaluation details. We evaluate performance using several metrics, including mean absolute relative error (Abs Rel), squared relative error (sq_rel), root mean square error (RMSE), root mean square error in log space ($\text{RMSE}_{\text{log}}$), absolute error in log space ($\text{log}_{\text{10}}$) and threshold accuracy ($\delta_{i}$).
Experimental Results
Quantitative results. We present the results on the NYUv2 and KITTI datasets in Table 1. (More detailed quantitative results are provided in Table 5 and Table 6 in the Supplementary Material.). Our approach consistently outperforms RSA(Zeng et al. 2024b) across all evaluation metrics and achieves performance comparable to scale recovery using ground-truth depths, as indicated in the Least Squares and Levenberg-Marquardt sections of the quantitative tables. The quantitative results show that models trained on a single dataset (VGLD-N or VGLD-K) perform slightly better within their respective domains compared to the unified model VGLD-NK. For example, VGLD-N/K-TCI with DAV2-ViTS as the RDE model achieves the best performance across all three evaluation metrics reported in the Table 1. Thanks to the precise routing capability of the DRM module, the performance gap between the single-dataset and unified models remains marginal, highlighting the strong cross-domain generalization ability of the unified VGLD-NK model. For example, based on DAV2-ViTS, the VGLD-N-TCI model achieves an AbsRel of 0.125 on NYUv2, and the VGLD-K-TCI model achieves 0.152 on KITTI. The unified VGLD-NK-TCI model obtains AbsRel scores of 0.127 and 0.153 on NYUv2 and KITTI, respectively, representing decreases of less than 1.58% and 0.65%.
Furthermore, models utilizing visual embeddings (VGLD-XX-I) consistently outperform those relying solely on textual embeddings (VGLD-XX-T), validating the effectiveness of visual cues for scale prediction over purely linguistic prompts. For example, based on DAV1-ViTS, the VGLD-NK-T model achieves AbsRel scores of 0.142 and 0.148 on NYUv2 and KITTI, respectively. In comparison, VGLD-NK-I achieves AbsRel scores of 0.114 and 0.142 on the same datasets, corresponding to improvements of 24.5% and 4.2%, respectively. Building on this, we combine both visual and textual embeddings (VGLD-XX-TCI), allowing visual features to guide the semantic alignment of textual inputs. This integration yields modest but meaningful improvements, thereby effectively addressing the challenge of visually grounded linguistic disambiguation.
Notably, the improvement of VGLD-XX-TCI over VGLD-XX-T is less pronounced on KITTI compared to NYUv2. We attribute this to the lower variance in outdoor scene descriptions in KITTI, whereas indoor scenes in NYUv2 exhibit much greater diversity—such as bathrooms, kitchens, classrooms… This higher variability in textual descriptions benefits the model by providing richer cues for more accurate estimation of scene-specific scaling parameters.
For completeness, the Supplementary Material presents more extensive quantitative results and qualitative comparisons, including those from the zero-shot evaluation setting.
Sensitivity to Variations in Linguistic Descriptions. A single image can be described using multiple textual expressions. To investigate how linguistic variation affects metric depth scale recovery, we evaluate the influence of different textual inputs on VGLD’s performance. Figures 3 and 4 present qualitative comparisons on NYU and KITTI under three distinct text prompts. We observe that while the RSA method—relying solely on textual descriptions—is highly sensitive to phrasing, VGLD demonstrates significantly greater robustness, consistently producing stable predictions for both scale and shift. This is most evident in the third image of Figure 3: RSA accurately recovers the depth when paired with Text-3 (whose prediction closely matches the ground truth), but exhibits substantial errors with Text-1 and Text-2. In contrast, VGLD achieves stable and accurate scale recovery across all three descriptions (Text-1 to Text-3). Moreover, VGLD often outperforms RSA across evaluation metrics, further highlighting its ability to provide reliable scalar estimations. The corresponding quantitative results are provided in the Supplementary Material, along with the three textual descriptions used for each image.
Ablation Study
Effect of the DRM. As shown in Table 2, we conduct ablation studies on the Domain Router Mechanism (DRM). The results demonstrate that incorporating the DRM consistently improves the overall performance of VGLD across all four backbone models and significantly enhances its cross-domain generalization capability. The ablation studies are conducted based on the VGLD-XX-TCI model.
Models Method NYU KITTI
[…]
MiDas-1 w/o DRM 0.128 0.415 0.757 0.136 3.636 0.855
with DRM 0.120 0.414 0.863 0.121 3.559 0.874
MiDas-2 w/o DRM 0.158 0.513 0.740 0.198 4.987 0.728
with DRM 0.151 0.513 0.780 0.185 4.804 0.737
DAV2-vits w/o DRM 0.135 0.459 0.798 0.163 4.060 0.767
with DRM 0.127 0.434 0.835 0.153 3.980 0.772
DAV1-vits w/o DRM 0.122 0.437 0.847 0.145 4.327 0.748
with DRM 0.112 0.392 0.883 0.136 4.008 0.816
Table 2: Performance comparison on NYU and KITTI datasets with and w/o(without) DRM. Best results are in bold.
Effect of the LM loss. To investigate the effect of different weights of LM loss $\mathcal{L}_{lm}$ on model training, we vary the value of $\beta$ in equation 5 and train the VGLD-NK-TCI model based on the DAV1-vits RDE backbone. Evaluation on both the NYUv2 and KITTI datasets shown in Table 3 that the model achieves the best performance when $\beta=0.1$. Compared to completely removing the $\beta$ term ($\beta=0$), the model achieves a 2.7% improvement in AbsRel on NYUv2 and a significantly larger gain of approximately 20.5% on KITTI. This demonstrates the effectiveness of the $\mathcal{L}_{\text{lm}}$ constraint, particularly in more open outdoor environments, where stronger guidance is needed to stabilize the training trajectory.
$\beta$ NYU KITTI
[…]
0 0.115 0.403 0.874 0.164 4.856 0.781
0.001 0.116 0.413 0.869 0.161 4.204 0.779
0.01 0.113 0.399 0.879 0.146 4.010 0.791
0.1 0.112 0.392 0.883 0.136 4.008 0.816
1 0.115 0.397 0.868 0.162 4.701 0.778
Table 3: Ablation on LM loss for NYUv2 and KITTI datasets. Best results are in bold, second best are underlined.
Computational Complexity. As shown in Table 4, we present a comparison of model parameters and inference times between VGLD and RSA to quantify the computational resources required. All evaluations were conducted on a single NVIDIA RTX 3090 (24GB). This experiment is conducted using the DAV1-vits RDE backbone. The results indicate that the scalar predictor in VGLD is more lightweight and efficient compared to that of RSA. However, VGLD additionally incorporates a CLIP image encoder, which introduces an extra 14ms of inference time compared to RSA. Despite this overhead, VGLD offers a favorable trade-off: it achieves a 32.1%(Ref. to Tabel 1) improvement in Abs Rel on NYUv2 with an inference time of just 14.08ms increases and a modest parameters, making it a practical and efficient choice.
Components RSA VGLD (ours)
Params# Inf. Times Params# Inf. Times
DAV1-vits 24.78M 9.62ms 24.78M 9.62ms
CLIP Text Encoder 63.43M 13.61ms 63.43M 13.61ms
CLIP Image Encoder - - 86.19M 14.90ms
Scalars Predictor 1.49M 1.76ms 1.18M 0.94ms
Total 89.7M 24.99ms 175.58M 39.07ms
Increase / M (ms) - - 85.88M $\uparrow$ 14.08ms $\uparrow$
Table 4: Computational Complexity Analysis. As shown in the table, the increase in model parameters(Params#) and inference times(Inf. Times) of VGLD compared to the RSA model primarily stems from the additional CLIP Image Encoder component.
Conclusion
We presented VGLD, a novel framework for monocular depth scale recovery that performs Visually-Guided Linguistic Disambiguation. VGLD leverages high-level visual semantics to resolve inconsistencies in textual inputs, enabling stable and accurate scale prediction across diverse linguistic descriptions. By jointly encoding image and text via CLIP and predicting global transformation parameters with an MLP, VGLD transforms relative depth maps into metric estimates in a robust and consistent manner. Extensive evaluations on both indoor and outdoor benchmarks show that VGLD significantly reduces estimation variance under different captions and generalizes well across domains. Empowered by a Domain Router Mechanism, VGLD further supports universal deployment across scene types. Compared to sensor-based methods, VGLD offers a lightweight and effective alternative for reliable scale alignment.
Limitations and future work.
Although linguistic-based scale recovery under visually-guided methods is highly robust, VGLD is still influenced by language modality. For different descriptions of the same image, the VGLD model may output inconsistent results (albeit with small error margins), especially when incorrect descriptions are used (e.g., describing an indoor scene as ”a photo of a narrow street.”). To address this issue, one feasible approach could be to further match the similarity between the language and image modalities, effectively excluding erroneous image descriptions. Future work could expand the image modality-assisted features of VGLD to enable more robust and fine-grained scale estimation, as well as enhance the model’s ability to handle malicious attacks in text descriptions.
Appendix A Supplementary Material
Evaluation Metrics
We evaluate our approach using the standard five error metrics and three accuracy metrics commonly adopted in prior works(Shao et al. 2023). Specifically, the error metrics include absolute mean relative error (Abs Rel), square relative error (sq_rel), log error($\log_{10}$), root mean squared error (RMSE), and its logarithmic variant ($\text{RMSE}_{\text{log}}$). The accuracy metrics are based on the percentage of inlier pixels ($\delta$) within three thresholds: $\delta_{1}<1.25$, $\delta_{2}<1.25^{2}$, and $\delta_{3}<1.25^{3}$.
  • Abs Rel: $\frac{1}{M}\sum_{(i,j)\in\Omega}|\hat{d}_{pred}(i,j)-d_{gt}(i,j)|/d_{gt}(i,j)$
  • sq_rel: $\frac{1}{M}\sum_{(i,j)\in\Omega}[(\hat{d}_{pred}(i,j)-d_{gt}(i,j))/d_{gt}(i,j)]^{2}$
  • RMSE: $\sqrt{\frac{1}{M}\sum_{(i,j)\in\Omega}(\hat{d}_{pred}(i,j)-d_{gt}(i,j))^{2}}$
  • $\text{RMSE}_{\text{log}}$: $\sqrt{\frac{1}{M}\sum_{(i,j)\in\Omega}(\log\hat{d}_{pred}(i,j)-\log d_{gt}(i,j))^{2}}$
  • $\text{log}_{10}$: $\frac{1}{M}\sum_{(i,j)\in\Omega}|\log_{10}(\hat{d}_{pred}(i,j))-\log_{10}(d_{gt}(i,j))|$
  • $\text{D}<thr$: $(max(\frac{\hat{d}_{pred}}{d_{gt}},\frac{d_{gt}}{\hat{d}_{pred}}))$ , $thr=1.25,1.25^{2},1.25^{3}$
Training details
The proposed VGLD is implemented in PyTorch2.0.1+CUDA11.8. We use the Adam optimizer with parameters $(\beta_{1},\beta_{2},\text{wd})=(0.9,0.999,0.001)$ and a learning rate of $3\times 10^{-4}$. All models are trained for 24 epochs on a single NVIDIA RTX 3090 GPU with 24GB of memory, , running in Ubuntu 22.04. The batch size is set to 6, and the total training time for each model is approximately 19 to 22 hours.
Qualitative comparisons
We present comparison examples of VGLD and baseline methods on the NYUv2 and KITTI datasets in Figure 5 and Figure 6, respectively. The error maps display the absolute relative error, where the overall brightness of the error maps clearly indicates the performance of our method. Notably, our approach achieves performance very close to that of the Levenberg-Marquardt fitting (LM Fit) across different scenes, demonstrating robust metric depth scale recovery. In contrast to the fixed scale and shift estimates produced by RSA, VGLD significantly improves the accuracy of depth predictions, with darker error maps indicating reduced error. Note: All qualitative comparison results in the VGLD section are inferred from the VGLD-NK-TCI method, where the RDE model used is DAV1-vits.
Quantitative Results on Sensitivity to Linguistic Description Variations
As shown in Table 7 and Table 9, We quantitatively evaluated the inference results and sensitivity of the VGLD model to variations in linguistic descriptions. For both indoor and outdoor datasets, three images were used, with each image paired with three distinct textual descriptions. The corresponding visualization figures are provided in Figure 3 and Figure 4(within the main text).. And the specific textual descriptions are provided in Table 8 and Table 10.
From the tables, it is evident that the VGLD model demonstrates greater robustness when processing three different textual descriptions, while the RSA model exhibits larger errors. Moreover, under identical textual descriptions, VGLD consistently outperforms RSA.
Zero-shot Generalization
Benefiting from the smaller domain gap of language descriptions across diverse scenes(Zeng et al. 2024a, b) and the ability of corresponding images to accurately indicate domain context, we conduct a zero-shot transfer experiment to demonstrate the generalization capability of VGLD. We evaluate the models on the SUN-RGBD(Song et al. 2015) , DIML Indoor(Cho et al. 2021), and DDAD(Guizilini et al. 2020) datasets without any fine-tuning. As shown in Figure 7, Figure 8, Figure 9 (qualitative results) and Table 11, Table 12, Table 13 (quantitative results), VGLD consistently outperforms baseline methods and produces results that closely match those fitted by the LM method. This demonstrates that, under visual guidance, VGLD maintains stable scalars estimation and exhibits enhanced generalization capabilities. Note that all zero-shot experiments are conducted using the VGLD-NK-TCI model built upon the DAV1-vits RDE backbone.
Effect of the initial seeds
To ensure the robustness of our training and verify that the results are not due to random initialization, we trained the model using three different random seeds. As illustrated in Figure 10, the resulting error bars indicate that variations due to different seeds are minimal, with nearly zero deviation.
Prompts for Natural Text Generation
To generate natural and semantically rich image descriptions—rather than relying on fixed prompt templates—we employ two vision-language models: LLaVA-v1.6-Vicuna-7B and LLaVA-v1.6-Mistral-7B(Jia et al. 2022). To ensure diversity in the generated captions, each model is prompted using six distinct instruction templates. These prompt templates are listed in Table 14.
Figure 5: Visualization of depth estimation on the NYUv2 dataset. The LM Fit represents the result obtained using the Levenberg-Marquardt algorithm. Note: Zeros in the ground truth indicate the absence of valid depth values (represented in black or deep red).
Figure 6: Visualization of depth estimation on the KITTI dataset. The LM Fit represents the result obtained using the Levenberg-Marquardt algorithm. Note: Zeros in the ground truth indicate the absence of valid depth values (represented in black or deep red).
Models Methods Abs Rel $\downarrow$ sq_rel $\downarrow$ RMSE $\downarrow$ $\text{RMSE}_{\text{log}}$downarrow log10 $\downarrow$ D1 $\uparrow$ D2 $\uparrow$ D3 $\uparrow$
ZoeDept(Bhat et al. 2023)h robust depth estimation 0.077 0.277 0.033 0.953 0.995 0.999
ZeroDepth(Guizilini et al. 2023) 0.074 0.269 0.103 0.954 0.995 1.000
Metric3Dv2(Hu et al. 2024a) 0.047 0.183 0.020 0.989 0.998 1.000
MiDas-1(Reiner et al. 2023) Least Squares 0.121 0.073 0.388 0.338 0.068 0.866 0.959 0.978
Levenberg Marquardt 0.056 0.021 0.218 0.080 0.024 0.969 0.995 0.998
RSA-N(Zeng et al. 2024b) 0.171 0.569 0.072 0.731 0.955 0.993
RSA-NK(Zeng et al. 2024b) 0.168 0.561 0.071 0.737 0.959 0.993
VGLD-N-T (Ours) 0.158 0.113 0.529 0.181 0.068 0.758 0.965 0.994
VGLD-N-I (Ours) 0.121 0.068 0.423 0.146 0.053 0.860 0.985 0.998
VGLD-N-TCI (Ours) 0.119 0.067 0.414 0.142 0.051 0.867 0.984 0.998
VGLD-NK-T (Ours) 0.159 0.113 0.526 0.178 0.067 0.751 0.971 0.995
VGLD-NK-I (Ours) 0.123 0.070 0.426 0.147 0.053 0.855 0.982 0.998
VGLD-NK-TCI (Ours) 0.120 0.068 0.414 0.143 0.052 0.863 0.984 0.998
MiDas-2(Ranftl et al. 2020) Least Squares 0.130 0.085 0.421 0.286 0.066 0.845 0.956 0.980
Levenberg Marquardt 0.094 0.049 0.330 0.122 0.039 0.916 0.985 0.997
VGLD-N-T (Ours) 0.180 0.140 0.596 0.212 0.078 0.688 0.946 0.990
VGLD-N-I (Ours) 0.154 0.106 0.524 0.186 0.067 0.775 0.960 0.993
VGLD-N-TCI (Ours) 0.151 0.104 0.507 0.181 0.064 0.789 0.964 0.993
VGLD-NK-T (Ours) 0.182 0.147 0.615 0.217 0.080 0.682 0.939 0.989
VGLD-NK-I (Ours) 0.155 0.108 0.520 0.185 0.066 0.776 0.961 0.992
VGLD-NK-TCI (Ours) 0.151 0.104 0.513 0.183 0.065 0.780 0.964 0.993
DAV2-vits(Yang et al. 2024b) Least Squares 0.122 0.074 0.392 0.362 0.070 0.866 0.959 0.977
Levenberg Marquardt 0.052 0.021 0.209 0.077 0.022 0.969 0.992 0.998
VGLD-N-T (Ours) 0.163 0.119 0.546 0.191 0.073 0.713 0.964 0.994
VGLD-N-I (Ours) 0.128 0.074 0.433 0.154 0.057 0.830 0.983 0.995
VGLD-N-TCI (Ours) 0.125 0.073 0.423 0.152 0.055 0.842 0.984 0.995
VGLD-NK-T (Ours) 0.161 0.115 0.539 0.189 0.073 0.714 0.967 0.994
VGLD-NK-I (Ours) 0.127 0.074 0.436 0.155 0.057 0.835 0.982 0.995
VGLD-NK-TCI (Ours) 0.127 0.074 0.434 0.155 0.057 0.835 0.981 0.995
DAV1-vits(Yang et al. 2024a) Least Squares 0.121 0.075 0.397 0.327 0.067 0.863 0.959 0.979
Levenberg Marquardt 0.057 0.022 0.230 0.081 0.024 0.967 0.995 0.999
RSA-N(Zeng et al. 2024b) 0.147 0.484 0.065 0.775 0.975 0.997
RSA-NK(Zeng et al. 2024b) 0.148 0.498 0.065 0.776 0.974 0.996
VGLD-N-T (Ours) 0.145 0.094 0.496 0.170 0.064 0.792 0.974 0.997
VGLD-N-I (Ours) 0.115 0.061 0.405 0.141 0.051 0.872 0.987 0.998
VGLD-N-TCI (Ours) 0.112 0.058 0.390 0.135 0.049 0.887 0.988 0.998
VGLD-NK-T (Ours) 0.142 0.089 0.483 0.168 0.063 0.787 0.979 0.997
VGLD-NK-I (Ours) 0.114 0.061 0.404 0.138 0.050 0.880 0.987 0.999
VGLD-NK-TCI (Ours) 0.112 0.059 0.392 0.135 0.048 0.883 0.988 0.999
Table 5: More detailed quantitative depth comparison on the NYUv2 dataset. Best results are in bold, second best are underlined.
Models Methods Abs Rel $\downarrow$ sq_rel $\downarrow$ RMSE $\downarrow$ $\text{RMSE}_{\text{log}}$downarrow log10 $\downarrow$ D1 $\uparrow$ D2 $\uparrow$ D3 $\uparrow$
ZoeDepth(Bhat et al. 2023) robust depth estimation 0.054 2.281 0.082 0.971 0.996 0.999
ZeroDepth(Guizilini et al. 2023) 0.053 2.087 0.083 0.968 0.995 0.999
Metric3Dv2(Hu et al. 2024a) 0.044 1.985 0.064 0.985 0.998 0.999
MiDas-1(Reiner et al. 2023) Least Squares 0.333 2.094 6.901 1.731 0.293 0.408 0.790 0.879
Levenberg Marquardt 0.091 0.425 3.373 0.127 0.038 0.925 0.987 0.996
RSA-K(Zeng et al. 2024b) 0.163 4.082 0.185 0.798 0.948 0.981
RSA-NK(Zeng et al. 2024b) 0.160 4.232 0.194 0.782 0.946 0.980
VGLD-K-T(Ours) 0.133 0.608 3.755 0.162 0.056 0.854 0.975 0.993
VGLD-K-I(Ours) 0.120 0.526 3.668 0.152 0.051 0.868 0.979 0.995
VGLD-K-TCI(Ours) 0.120 0.523 3.598 0.151 0.051 0.871 0.980 0.996
VGLD-NK-T(Ours) 0.130 0.568 3.744 0.161 0.056 0.844 0.975 0.995
VGLD-NK-I(Ours) 0.122 0.543 3.574 0.151 0.051 0.868 0.979 0.995
VGLD-NK-TCI(Ours) 0.120 0.528 3.559 0.150 0.051 0.874 0.980 0.996
MiDas-2(Ranftl et al. 2020) Least Squares 0.336 2.172 6.925 1.658 0.283 0.421 0.778 0.876
Levenberg Marquardt 0.155 0.770 4.190 0.185 0.062 0.809 0.966 0.990
VGLD-K-T(Ours) 0.194 1.290 5.030 0.225 0.079 0.709 0.930 0.981
VGLD-K-I(Ours) 0.183 1.154 4.842 0.215 0.075 0.733 0.942 0.983
VGLD-K-TCI(Ours) 0.178 1.146 4.806 0.210 0.073 0.748 0.942 0.984
VGLD-NK-T(Ours) 0.191 1.260 4.994 0.221 0.078 0.723 0.932 0.981
VGLD-NK-I(Ours) 0.184 1.179 4.808 0.213 0.074 0.740 0.938 0.984
VGLD-NK-TCI(Ours) 0.180 1.158 4.804 0.212 0.074 0.737 0.943 0.984
DAV2-vits(Yang et al. 2024b) Least Squares 0.330 2.053 6.737 1.729 0.292 0.423 0.790 0.877
Levenberg Marquardt 0.103 0.454 3.277 0.135 0.042 0.919 0.987 0.997
VGLD-K-T(Ours) 0.166 0.822 4.189 0.190 0.070 0.756 0.953 0.992
VGLD-K-I(Ours) 0.154 0.698 4.219 0.187 0.067 0.756 0.966 0.995
VGLD-K-TCI(Ours) 0.152 0.657 3.872 0.179 0.065 0.779 0.972 0.996
VGLD-NK-T(Ours) 0.164 0.786 4.287 0.193 0.070 0.752 0.955 0.993
VGLD-NK-I(Ours) 0.160 0.748 4.031 0.187 0.069 0.761 0.965 0.995
VGLD-NK-TCI(Ours) 0.153 0.695 3.980 0.180 0.066 0.772 0.973 0.996
DAV1-vits(Yang et al. 2024a) Least Squares 0.331 2.078 6.772 1.714 0.291 0.423 0.786 0.875
Levenberg Marquardt 0.112 0.495 3.375 0.142 0.045 0.897 0.986 0.997
RSA-K(Zeng et al. 2024b) 0.160 4.437 0.189 0.780 0.958 0.988
RSA-NK(Zeng et al. 2024b) 0.158 4.457 0.179 0.786 0.967 0.987
VGLD-K-T(Ours) 0.151 0.747 4.354 0.186 0.066 0.773 0.963 0.994
VGLD-K-I(Ours) 0.144 0.646 4.074 0.178 0.063 0.790 0.975 0.996
VGLD-K-TCI(Ours) 0.140 0.686 4.081 0.172 0.061 0.807 0.975 0.996
VGLD-NK-T(Ours) 0.148 0.728 4.293 0.183 0.065 0.781 0.966 0.995
VGLD-NK-I(Ours) 0.142 0.759 4.151 0.172 0.061 0.814 0.975 0.996
VGLD-NK-TCI(Ours) 0.136 0.632 4.008 0.169 0.059 0.816 0.977 0.997
Table 6: More detailed quantitative depth comparison on the KITTI dataset. Best results are in bold, second best are underlined.
Idx Text-idx Method Abs Rel $\downarrow$ RMSE $\downarrow$ D1 $\uparrow$ pred_shift LM_shift pred_scale LM_scale
1 Text-1 RSA(Zeng et al. 2024b) 0.210 0.689 0.263 1.255 1.032
VGLD 0.080 0.240 0.987 1.207 1.020
Text-2 RSA(Zeng et al. 2024b) 0.110 0.367 0.995 1.220 1.028
VGLD 0.065 0.271 0.999 1.220 1.284
Text-3 RSA(Zeng et al. 2024b) 0.054 0.216 0.997 1.210 1.026
VGLD 0.052 0.232 0.997 1.216 1.193 1.025 1.026
2 Text-1 RSA(Zeng et al. 2024b) 0.089 0.344 0.961 1.202 1.032
VGLD 0.073 0.271 0.962 1.185 1.034
Text-2 RSA(Zeng et al. 2024b) 0.069 0.256 0.956 1.214 1.030
VGLD 0.063 0.272 0.962 1.213 1.037
Text-3 RSA(Zeng et al. 2024b) 0.064 0.296 0.947 1.218 1.036
VGLD 0.062 0.280 0.954 1.228 1.210 1.034 1.033
3 Text-1 RSA(Zeng et al. 2024b) 0.251 0.868 0.138 1.331 1.042
VGLD 0.147 0.433 0.920 1.254 1.041
Text-2 RSA(Zeng et al. 2024b) 0.055 0.240 0.993 1.199 1.035
VGLD 0.054 0.170 0.994 1.205 1.038
Text-3 RSA(Zeng et al. 2024b) 0.058 0.154 0.994 1.212 1.039
VGLD 0.055 0.138 0.994 1.199 1.218 1.035 1.034
Table 7: Quantitative results on the NYUv2 dataset comparing VGLD and RSA in response to different textual descriptions. The LM_shift and LM_scale represent scalars values fitted using the Levenberg-Marquardt method. Best results are in bold.
Idx Texts-idx Text Description
1 Text-1 A man is standing in a doorway, looking at a bed with a striped comforter.
Text-2 The bed is positioned in the corner of the room, with a man standing in the doorway, and a fish tank nearby.
Text-3 A man stands in a doorway, looking into a bedroom with a large bed, a wooden dresser, and a fish tank.
2 Text-1 The image shows a classroom with a play area, a table with chairs, and a sink.
Text-2 The image shows a classroom with a table, chairs, and a sink, all situated near a wall with bulletin boards and a window.
Text-3 The image shows a classroom with a table, chairs, a sink, a bulletin board, a bookshelf, a window, and a rug.
3 Text-1 The image shows a red couch with towels hanging over the back, a flat screen television, and a framed jersey on the wall.
Text-2 The image shows a red couch with a pink towel and a blue towel on it, positioned in front of a television with a framed jersey on the wall behind it.
Text-3 The image shows a living room with a red couch, a flat screen TV, a framed jersey, and a guitar.
Table 8: The table shows three distinct textual descriptions provided for each image in the NYUv2 dataset, used as linguistic inputs for evaluating model sensitivity.
Idx Text-idx Method Abs Rel $\downarrow$ RMSE $\downarrow$ D1 $\uparrow$ pred_shift LM_shift pred_scale LM_scale
1 Text-1 RSA(Zeng et al. 2024b) 0.097 4.060 0.926 1.005 1.011
VGLD 0.075 3.562 0.949 1.004 1.001
Text-2 RSA(Zeng et al. 2024b) 0.084 4.251 0.923 1.007 1.010
VGLD 0.077 3.067 0.949 1.004 1.010
Text-3 RSA(Zeng et al. 2024b) 0.072 3.140 0.952 1.004 1.010
VGLD 0.068 3.088 0.951 1.004 1.003 1.010 1.010
2 Text-1 RSA(Zeng et al. 2024b) 0.108 2.327 0.905 1.009 1.014
VGLD 0.063 1.887 0.984 1.135 1.015
Text-2 RSA(Zeng et al. 2024b) 0.281 5.341 0.537 1.006 1.013
VGLD 0.099 2.144 0.915 1.010 1.014
Text-3 RSA(Zeng et al. 2024b) 0.109 2.157 0.906 1.011 1.014
VGLD 0.073 1.861 0.952 1.011 1.009 1.015 1.017
3 Text-1 RSA(Zeng et al. 2024b) 0.268 6.526 0.740 1.004 1.009
VGLD 0.119 2.516 0.919 1.009 1.010
Text-2 RSA(Zeng et al. 2024b) 0.171 4.479 0.854 1.005 1.010
VGLD 0.077 2.287 0.942 1.008 1.010
Text-3 RSA(Zeng et al. 2024b) 0.081 2.421 0.938 1.007 1.010
VGLD 0.062 2.236 0.953 1.009 1.008 1.011 1.011
Table 9: Quantitative results on the KITTI dataset comparing VGLD and RSA in response to different textual descriptions. The LM_shift and LM_scale represent scalars values fitted using the Levenberg-Marquardt method. Best results are in bold.
Idx Text-idx Text Description
1 Text-1 The image shows a narrow city street lined with parked cars and buildings on both sides.
Text-2 The image shows a narrow street lined with parked cars and buildings, with a clear sky overhead.
Text-3 The image shows a narrow street with parked cars on both sides, leading towards a building with a red awning.
2 Text-1 The image shows a narrow alleyway with a white gate at the end, a bridge overhead, and a hillside on one side.
Text-2 The image shows a narrow alleyway with a white gate, a fence, a building, a bridge, and a sign, all situated in close proximity to each other.
Text-3 A narrow alleyway with a white gate, a fence, a building, a bridge, a tree, a sign, and a hill.
3 Text-1 The image captures a lively street scene with people walking and riding bicycles, shops and buildings lining the street, and a clear blue sky overhead.
Text-2 The image shows a narrow street in a European city, with buildings on both sides, a pedestrian walkway in the middle, and people walking and biking on the street.
Text-3 The image shows a bustling city street with people walking and riding bicycles, shops and buildings lining the street, and a clear blue sky overhead.
Table 10: The table shows three distinct textual descriptions provided for each image in the KITTI dataset, used as linguistic inputs for evaluating model sensitivity.
Figure 7: Zero-shot generalization on the SUN-RGBD dataset(Indoor). The models are evaluated without any fine-tuning. Benefiting from robust scale prediction, our VGLD method produces depth maps that are significantly closer to the ground truth compared to RSA.
RDE Model Method Lower is better Higher is better
Abs Rel $\downarrow$ sq_rel $\downarrow$ RMSE $\downarrow$ $\text{RMSE}_{\text{log}}$downarrow log10 $\downarrow$ D1 $\uparrow$ D2 $\uparrow$ D3 $\uparrow$
ZoeDepth(Bhat et al. 2023) robust depth $\dagger$ 0.123 0.356 0.053 0.856 0.979 0.995
ScaleDepth(Zhu et al. 2024) 0.129 0.359 0.866
MiDas-1(Reiner et al. 2023) Least Squares 0.197 0.418 0.346 0.278 0.061 0.873 0.964 0.981
Levenberg Marquardt 0.158 0.440 0.252 0.116 0.032 0.950 0.988 0.995
RSA-NK(Zeng et al. 2024b) 0.299 0.589 0.575 0.251 0.094 0.615 0.900 0.977
VGLD-NK-T(Ours) 0.318 0.647 0.566 0.242 0.089 0.643 0.914 0.980
VGLD-NK-I(Ours) 0.259 0.595 0.468 0.202 0.071 0.751 0.957 0.991
VGLD-NK-TCI(Ours) 0.262 0.628 0.467 0.202 0.071 0.751 0.959 0.991
MiDas-2(Ranftl et al. 2020) Least Squares 0.203 0.419 0.365 0.272 0.062 0.860 0.961 0.981
Levenberg Marquardt 0.173 0.438 0.291 0.132 0.039 0.926 0.984 0.994
VGLD-NK-T(Ours) 0.316 0.795 0.597 0.249 0.090 0.639 0.908 0.978
VGLD-NK-I(Ours) 0.288 0.688 0.552 0.246 0.090 0.627 0.922 0.984
VGLD-NK-TCI(Ours) 0.275 0.670 0.513 0.225 0.080 0.694 0.941 0.987
DAV2-vits(Yang et al. 2024b) Least Squares 0.194 0.418 0.337 0.305 0.062 0.880 0.963 0.980
Levenberg Marquardt 0.146 0.439 0.224 0.103 0.027 0.961 0.989 0.995
VGLD-NK-T(Ours) 0.304 0.742 0.564 0.236 0.089 0.644 0.920 0.983
VGLD-NK-I(Ours) 0.273 0.564 0.535 0.236 0.090 0.617 0.931 0.989
VGLD-NK-TCI(Ours) 0.241 0.545 0.433 0.189 0.067 0.779 0.967 0.993
DAV1-vits(Yang et al. 2024a) Least Squares 0.196 0.416 0.341 0.282 0.061 0.875 0.963 0.981
Levenberg Marquardt 0.151 0.440 0.234 0.108 0.029 0.957 0.989 0.995
RSA-NK(Zeng et al. 2024b) 0.290 0.563 0.571 0.250 0.092 0.640 0.899 0.969
VGLD-NK-T(Ours) 0.281 0.583 0.532 0.214 0.078 0.711 0.945 0.987
VGLD-NK-I(Ours) 0.250 0.573 0.443 0.194 0.070 0.764 0.965 0.991
VGLD-NK-TCI(Ours) 0.241 0.545 0.433 0.189 0.067 0.779 0.967 0.993
Table 11: Zero-shot generalization to SUN-RGBD (Indoor). Best results are in bold, second best are underlined.
Figure 8: Zero-shot generalization on the DIML Indoor dataset(Indoor). The models are evaluated without any fine-tuning. Benefiting from robust scale prediction, our VGLD method produces depth maps that are significantly closer to the ground truth compared to RSA.
RDE Model Method Lower is better Higher is better
Abs Rel $\downarrow$ sq_rel $\downarrow$ RMSE $\downarrow$ $\text{RMSE}_{\text{log}}$downarrow log10 $\downarrow$ D1 $\uparrow$ D2 $\uparrow$ D3 $\uparrow$
MiDas-1(Reiner et al. 2023) Least Squares 0.123 0.070 0.364 0.357 0.069 0.868 0.959 0.978
Levenberg Marquardt 0.070 0.029 0.241 0.095 0.029 0.952 0.991 0.998
RSA-NK(Zeng et al. 2024b) 0.219 0.218 0.667 0.246 0.096 0.612 0.882 0.964
VGLD-NK-T (Ours) 0.251 0.385 0.683 0.240 0.094 0.622 0.898 0.969
VGLD-NK-I (Ours) 0.188 0.138 0.544 0.208 0.079 0.696 0.943 0.982
VGLD-NK-TCI (Ours) 0.212 0.281 0.623 0.228 0.088 0.638 0.930 0.978
MiDas-2(Ranftl et al. 2020) Least Squares 0.133 0.080 0.394 0.345 0.071 0.846 0.954 0.977
Levenberg Marquardt 0.086 0.039 0.285 0.114 0.036 0.929 0.988 0.996
VGLD-NK-T (Ours) 0.243 0.359 0.737 0.264 0.100 0.585 0.877 0.964
VGLD-NK-I (Ours) 0.235 0.201 0.722 0.294 0.115 0.460 0.849 0.975
VGLD-NK-TCI (Ours) 0.227 0.371 0.690 0.262 0.100 0.570 0.894 0.979
DAV2-vits(Yang et al. 2024b) Least Squares 0.123 0.068 0.361 0.361 0.069 0.872 0.960 0.978
Levenberg Marquardt 0.066 0.024 0.226 0.092 0.028 0.958 0.993 0.998
VGLD-NK-T (Ours) 0.228 0.300 0.673 0.246 0.096 0.593 0.891 0.981
VGLD-NK-I (Ours) 0.212 0.169 0.663 0.259 0.103 0.514 0.899 0.989
VGLD-NK-TCI (Ours) 0.196 0.487 0.610 0.208 0.082 0.678 0.952 0.990
DAV1-vits(Yang et al. 2024a) Least Squares 0.118 0.063 0.345 0.344 0.066 0.875 0.961 0.979
Levenberg Marquardt 0.056 0.020 0.203 0.081 0.024 0.970 0.994 0.999
RSA-NK(Zeng et al. 2024b) 0.216 0.283 0.679 0.249 0.098 0.608 0.873 0.964
VGLD-NK-T (Ours) 0.211 0.711 0.627 0.215 0.084 0.683 0.927 0.983
VGLD-NK-I (Ours) 0.193 0.200 0.597 0.220 0.087 0.619 0.950 0.994
VGLD-NK-TCI (Ours) 0.196 0.487 0.610 0.208 0.082 0.678 0.952 0.990
Table 12: Zero-shot generalization to DIML Indoor. Best results are in bold, second best are underlined.
Figure 9: Zero-shot generalization on the DDAD dataset(Outdoor). The models are evaluated without any fine-tuning. Benefiting from robust scale prediction, our VGLD method produces depth maps that are significantly closer to the ground truth compared to RSA. Note that due to the sparse ground truth depth maps in the DDAD dataset, the visualization quality is poor. Therefore, LM Fit is used as a substitute for the ground truth depth map in the visualizations.
RDE Model Method Lower is better Higher is better
Abs Rel $\downarrow$ sq_rel $\downarrow$ RMSE $\downarrow$ $\text{RMSE}_{\text{log}}$downarrow log10 $\downarrow$ D1 $\uparrow$ D2 $\uparrow$ D3 $\uparrow$
MiDas-1(Reiner et al. 2023) Least Squares 0.319 2.265 7.252 1.936 0.301 0.409 0.844 0.920
Levenberg Marquardt 0.201 1.231 5.411 0.223 0.079 0.673 0.960 0.991
RSA-NK(Zeng et al. 2024b) 0.223 - 19.342 0.325 - 0.631 0.903 0.966
VGLD-NK-T (Ours) 0.215 2.519 10.467 0.320 0.102 0.630 0.851 0.935
VGLD-NK-I (Ours) 0.212 2.409 10.061 0.311 0.101 0.633 0.851 0.935
VGLD-NK-TCI (Ours) 0.209 2.517 10.446 0.319 0.100 0.659 0.862 0.941
MiDas-2(Ranftl et al. 2020) Least Squares 0.328 2.447 7.490 1.902 0.298 0.407 0.828 0.914
Levenberg Marquardt 0.232 1.557 5.985 0.253 0.090 0.609 0.934 0.985
VGLD-NK-T (Ours) 0.232 2.625 14.324 0.326 0.112 0.603 0.841 0.936
VGLD-NK-I (Ours) 0.220 2.526 12.235 0.321 0.106 0.642 0.865 0.947
VGLD-NK-TCI (Ours) 0.212 2.521 10.032 0.311 0.102 0.659 0.881 0.954
DAV2-vits(Yang et al. 2024b) Least Squares 0.318 2.239 7.205 1.937 0.300 0.410 0.847 0.920
Levenberg Marquardt 0.173 1.027 4.988 0.200 0.069 0.757 0.974 0.992
VGLD-NK-T (Ours) 0.221 3.125 8.769 0.252 0.085 0.675 0.927 0.977
VGLD-NK-I (Ours) 0.185 2.848 8.344 0.232 0.074 0.746 0.929 0.980
VGLD-NK-TCI (Ours) 0.176 2.002 7.925 0.238 0.075 0.748 0.942 0.981
DAV1-vits(Yang et al. 2024a) Least Squares 0.316 2.223 7.182 1.932 0.299 0.411 0.850 0.920
Levenberg Marquardt 0.156 0.929 4.766 0.185 0.062 0.817 0.977 0.991
RSA-NK(Zeng et al. 2024b) 0.207 - 19.715 0.303 - 0.642 0.903 0.976
VGLD-NK-T (Ours) 0.210 2.598 13.432 0.318 0.108 0.708 0.913 0.970
VGLD-NK-I (Ours) 0.192 2.557 9.275 0.258 0.081 0.732 0.922 0.975
VGLD-NK-TCI (Ours) 0.186 2.403 8.984 0.246 0.079 0.742 0.932 0.975
Table 13: Zero-shot generalization to DDAD (Outdoor). Best results are in bold, second best are underlined.
Figure 10: Error bars showing performance variations across different random seeds (0, 1, 2) for Abs Rel, RMSE, and D1 metrics. Each group of bars corresponds to a specific variant of the VGLD model.
Idx Prompts
1 Summarize the image in one sentence.
2 Summarize the image in one sentence, focusing mainly on the proximity relationships of the objects.
3 Describe the image in one sentence from near to far, focusing on the absolute positions of objects, with no more than 8 categories.
4 Describe the image in one sentence from near to far, focusing on the objects’ relative positions, with no more than 8 categories.
5 Summarize the image in one sentence, describing the overall spatial layout of the image.
6 Summarize the image in one sentence, describing the overall distance relationships in the image.
Table 14: Prompts for Natural Text Generation. We utilize two LLaVA models(llava-v1.6-vicuna-7b and llava-v1.6-mistral-7b), each generating 6 textual descriptions per image, resulting in a total of 12 diverse descriptions for each image.